What is the mean and standard deviation of this aggregate data set?

18-23 13
24-30 15
31-40 6
41-50 3
51-60 6
61+ 1
Answer #1

First, We need to find out Midpoint:

Midpoint = (Upper class Limit + Lower class limit) / 2

Midpoint = (18+23)/2 = 20.5 (For, first class)

(24+30)/2 = 27 (For, second class)

and so on ...

Frequency Distribution Table:

Class interval Frequency (fi) Midpoints (Xi) Xi*Fi Fi*Xi2
18-23 13 20.5 266.5 5463.25
24-30 15 27 405 10935
31-40 6 35.5 213 7561.5
41-50 3 45.5 136.5 6210.75
51-60 6 55.5 333 18481.5
61+ 1 65.5 65.5 4290.25
Total 44 1419.5 52942.25

Mean:

1419.5 32.26 FiXi 44

Where,

N =fi

Standard Deviation:

(EFiXi2 (1419.5)2 Σ FiX? 52942.25 SD(S) N 44 12.89 N 1 44 1